- Approach and show your work in exactly the same way as demonstrated in the example below -   

Use the Master Theorem to characterize and solve the following recurrence equations by stating at the end which case was used and why:

  1. T(n) = 25T(n/5) + n

  1. T(n) = 36T(n/6) + (n log n)2

  1. T(n) = 8T(n/3) + n2

Theorem

       T(n) = c                          if n = 1

       T(n) = a T(n/b) + f(n)     if n > 1

Let f(n) be as above, where f(n) Θ(nk), k >= 0

  • Case 1: If a > bk           then       T(n) = Θ(n log b a)
  • Case 2: If a = bk           then        T(n) = Θ (nlog b a logk+1 n )
  • Case 3: If a < bk           then        T(n) = Θ(f(n) )

Case 1 Example:

T(n) = 2T(n/2) + 1

In this case, what is f(n)?
And what are the values for a, b, and k? (a =? b =? k =?)

f(n) = 1 (Because in terms of n, 1 = n0)

a = 2, b = 2, k = 0

Which of the cases applies?

a = 2, and bk = 20= 1

Since 2 > 1, a > bk=> case 1

Therefore, T(n) is Θ (n log 2 2) by the master theorem.

And since log 2 2= 1, that means T(n) є Θ(n).
Answer #1
  1. T(n) = 25T(n/5) + n

f(n) = n

a=25 , b=5 , k=1

25 > 51 so  a > bk [case 1 applies]

T(n) = Θ(n logba) = Θ(nlog525) = Θ(n2)

============================================================================================

2. T(n) = 36T(n/6) + (n log n)2

f(n) = (n log n)2

a=36 , b=6 , k=2

36 == 62 so  a == bk [case 2 applies]

T(n) =T(n) = Θ (nlogba logk+1 n ) = Θ(nlog636 log3n) = Θ(n2 log3n)

============================================================================================

  1. T(n) = 8T(n/3) + n2

f(n) = n2

a=8 , b=3 , k=2

8 < 32 so  a < bk [case 3 applies]

T(n) = Θ(f(n)) = Θ(n2)

============================================================================================